OpenStudy (toxicsugar22):

coverics at (3,7) and (3,-1) major axis of length 10 and can u show me how u did this

OpenStudy (amistre64):

coverics?

OpenStudy (toxicsugar22):

covertices

OpenStudy (amistre64):

vertices; is this ellipse or hyperbola?

OpenStudy (toxicsugar22):

ellipse

OpenStudy (amistre64):

cant really determine an ellipse from just this info.... at best I can do is: (x-3)^2 (y-3)^2 ------ + ------ = 1 b^2 16

OpenStudy (amistre64):

without knowing a focus; or eccentricity; there is no solid way to determine the 'b' value

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

i got it

OpenStudy (anonymous):

yes there are infinitely many ellipses with major axis 10 and those vertices

OpenStudy (anonymous):

wait nevermind that

OpenStudy (anonymous):

there's only one it has minor axis 6

OpenStudy (toxicsugar22):

vertics at (-4,9) and (-4,-3), Covertices at (-7,3) and (-1,3)

OpenStudy (toxicsugar22):

and cam u show me how u got that

OpenStudy (toxicsugar22):

thant is a question

OpenStudy (toxicsugar22):

can u help me with that

OpenStudy (anonymous):

you look at the distance between vertices to find the axis lengths

OpenStudy (anonymous):

-7-1=6 and 9-(-3)=12

OpenStudy (anonymous):

the center is (-4,3)

OpenStudy (anonymous):

you can get your equation from that. (x+4)^2/(9)+(y-3)^2/(36)

OpenStudy (anonymous):

=1

OpenStudy (toxicsugar22):

8x^2+y^2-48+4y+68=0

OpenStudy (toxicsugar22):

write the equation in standard form